# RNAGenScape: Property-guided Optimization and Interpolation of mRNA Sequences with Manifold Langevin Dynamics

### v1

@@ 4 Empirical Results
In this section, we demonstrate the effectiveness of RNAGenScape on two key tasks: (1) mRNA sequence optimization and (2) mRNA sequence interpolation. The first task is broadly relevant to applications in therapeutics and synthetic biology. For example, enhancing the translation efficiency and stability of an mRNA vaccine can increase its protein yield and persistence, thereby boosting therapeutic efficacy while reducing the required dose. The second task facilitates the exploration of intermediate variants. This can provide insights into the functional landscape of regulatory elements within mRNAs of interest.
@@ 4.1 Experimental Settings
Datasets and tasks
We evaluate RNAGenScape on three mRNA datasets that capture diverse contexts and experimental designs. The optimization objectives are underlined.
@@
    Zebrafish includes five subsets of zebrafish 5’ UTR, experimentally measured using nascent protein-transducing ribosome affinity purification (Strayer et al., 2023), a massively parallel reporter assay to quantify translation control. It spans various stages and conditions of development, and each subset contains approximately 11,000 5’ UTR sequences each with 124 nucleotides along with annotations on translation efficiency.
@@
    OpenVaccine contains 2,400 mRNA sequences devised for COVID-19 mRNA vaccines, each with 107 nucleotides (Das et al., 2020). They are collected with degradation profiles under multiple conditions, quantified to mRNA stability relevant to vaccine design.
@@
    Ribosome-loading is a large-scale library of approximately 260,000 5’ UTR sequences each with 50 nucleotides (Sample et al., 2019), paired with pseudouridine-modified coding sequences of enhanced green fluorescent protein. The sequences are annotated on mean ribosome load, a property that reflects translation efficiency.
Figure 2: Latent space trajectories of RNAGenScape over 10 optimization steps. The trajectories follow smooth and reasonable paths with steady improvement in the target property. (a) Trajectories in the PHATE space. (b) 2D structures. (c) 3D structures. (d) Intermediate products of various methods midway through optimization (5th step).
Baselines
We compared our method with a range of popular de novo generative modeling approaches, including variational autoencoder (VAE) (Kingma et al., 2013), Wasserstein generative adversarial network with gradient penalty regularization (WGAN-GP) (Gulrajani et al., 2017), denoising diffusion probabilistic model (DDPM) (Ho et al., 2020), latent diffusion model (LDM) (Rombach et al., 2022), and flow matching (FM) (Rombach et al., 2022). We also included classic optimization methods, including gradient ascent (Williams, 1992; Zinkevich, 2003), Markov chain Monte Carlo (MCMC) (Brooks, 1998; Andrieu et al., 2003), and hill climbing (Selman & Gomes, 2006). All classic optimization baselines were GPU-compatible adaptions from the implementation in (Castro et al., 2022). Lastly, we benchmarked against optimization methods originally designed for proteins, DiffAb (Luo et al., 2022), IgLM (Shuai et al., 2023), and NOS (Gruver et al., 2023).
Evaluation
Since the optimization process could and should result in mRNA sequences not covered by the dataset, to quantify their properties, we trained a separate property prediction model $\mathcal{P}_{\text{oracle}}(x)$ to serve as a proxy of the ground truth. $\mathcal{P}_{\text{oracle}}(x)$ is used for evaluation only, and is strictly invisible during inference to avoid circular dependency.
@@
All experiments were performed under 5 random seeds and the average results are reported. Hyperparameters and hardware used are summarized in Appendix A.
4.2 RNAGenScape produces structured, data-aligned trajectories
RNAGenScape operates within a learned latent space that reflects the manifold of real biological sequences. To illustrate this behavior, we visualize individual optimization runs in Figure 2. The trajectory exhibits monotonic increases in the target property, while remaining near regions populated by real sequences. See Appendix D for more examples. These trajectories are direct consequences of the manifold-constrained dynamics, which guided each step toward high-property regions while staying on the manifold.
Importantly, all intermediate steps during optimization can be decoded into mRNA sequences, allowing researchers to examine how sequences evolve step by step as specific properties are optimized. As a qualitative illustration of biological plausibility, we show that intermediate results can be properly folded by ViennaRNA (Lorenz et al., 2011) and RhoFold (Shen et al., 2024) into 2D and 3D structures (Figure 2b-c). In contrast, folding intermediate products of several other methods lead to failure, as indicated by the broken structures (Figure 2d).
Table 1: Our proposed RNAGenScape achieves superior property optimization while also being inference-efficient. Top performers among property optimization methods are bolded. For de novo generative models, the optimization columns are grayed out, as they cannot explicitly steer properties; reported values instead reflect their learned distributions. $\Delta$ denotes the median change in property. $\%$ denotes success rate (the percentage of mRNAs improved).
Methods
@@
property = mean ribosome load
Inference Speed ++property -property ++property -property ++property -property
ms/sample $\downarrow$ $\Delta$ $\uparrow$ % $\uparrow$ $\Delta$ $\downarrow$ % $\uparrow$ $\Delta$ $\uparrow$ % $\uparrow$ $\Delta$ $\downarrow$ % $\uparrow$ $\Delta$ $\uparrow$ % $\uparrow$ $\Delta$ $\downarrow$ % $\uparrow$
@@
VAE (Kingma et al., 2013) 000.13 -0.80 67.7 -0.80 32.3 -0.41 40.0 -0.41 60.0 -0.37 38.3 -0.37 61.7
WGAN-GP (Gulrajani et al., 2017) 000.07 -1.23 16.4 -1.23 83.6 -0.02 47.9 -0.02 53.1 -1.16 22.5 -1.16 77.5
DDPM (Ho et al., 2020) 000.91 -0.16 55.4 -0.16 44.6 -0.28 64.6 -0.28 35.4 -0.28 40.4 -0.28 59.6
LDM (Rombach et al., 2022) 000.74 -0.43 36.5 -0.43 63.5 -1.38 78.0 -1.38 21.0 -1.11 46.0 -1.11 54.0
FM (Lipman et al., 2022) 005.82 -0.17 55.6 -0.17 44.4 -0.20 62.9 -0.20 37.1 -0.25 41.9 -0.25 58.1
@@
DiffAb (Luo et al., 2022) 041.04 -0.20 62.4 -0.17 41.1 -0.37 73.8 -0.40 26.2 -0.10 41.5 -0.16 60.8
IgLM (Shuai et al., 2023) 157.57 -0.07 52.9 -0.01 49.3 -0.07 54.8 -0.06 46.7 -0.42 69.6 -1.28 80.6
NOS-C (Gruver et al., 2023) 000.99 -0.03 48.6 -0.66 70.4 -0.96 90.0 -0.05 52.1 -0.21 42.6 -0.26 59.0
NOS-D (Gruver et al., 2023) 000.96 -0.22 57.1 -0.20 42.7 -0.46 71.8 -0.25 36.2 -0.25 41.3 -0.26 59.3
Sequence-space MCMC 003.84 -0.53 33.5 -0.54 67.1 -0.13 41.8 -0.18 58.7 -1.02 25.0 -1.56 83.8
OAE + Gradient Ascent 000.50 -0.51 33.9 -0.44 63.8 -0.31 66.3 -0.19 61.3 -0.47 34.7 -1.60 84.0
OAE + MCMC 010.93 -0.43 35.8 -0.44 64.2 -0.45 30.2 -0.17 60.0 -1.41 18.7 -1.42 81.3
OAE + Hill Climbing 081.52 -0.52 33.6 -0.56 67.5 -0.16 40.8 -0.16 60.0 -1.38 19.0 -1.39 81.0
OAE + Stochastic Hill Climbing 099.66 -0.51 33.5 -0.56 67.9 -0.12 43.5 -0.15 60.0 -1.40 20.0 -1.41 80.4
RNAGenScape without SUGAR (ours) [^1] The optimal SUGAR upsampling ratio is 0 for Ribosome-loading, and hence we have identical performance with and without SUGAR in that dataset. See ablation studies (Section 4.6). 000.57 -1.19 89.3 -0.83 74.2 -1.39 95.2 -0.33 65.2 -0.46 72.4 -1.66 87.3
RNAGenScape (ours) 000.57 -1.48 94.0 -1.32 85.6 -1.33 93.5 -0.87 86.9 -0.46 72.4 -1.66 87.3
4.3 RNAGenScape achieves superior property optimization
We quantitatively compare RNAGenScape against a range of de novo generative models and optimization baselines (Table 1). Although de novo approaches are effective in modeling the data distribution, they offer limited to no control over the target properties. As a result, their performance in property optimization is unfavorable.
Among property optimization methods, RNAGenScape consistently delivers the strongest results, achieving the highest median property change and the highest success rate in both the positive and negative directions. In particular, its median improvement is nearly twice that of the next-best method in several cases, and its success rate exceeds all other approaches.
Table 2: Manifold fidelity, represented by the average $\ell_{2}$ distance between the generated or optimized sequences to the data manifold. Results are averaged over all datasets.
Methods
latent space
distance $\downarrow$
VAE 0.737
WGAN-GP 1.729
DDPM 0.254
LDM 0.584
FM 0.259
DiffAb 0.237
IgLM 0.740
NOS-C 0.539
NOS-D 0.260
Sequence-space MCMC 0.292
OAE + Gradient Ascent 0.460
OAE + MCMC 0.297
OAE + Hill Climbing 0.300
OAE + Stochastic Hill Climbing 0.288
RNAGenScape without SUGAR (ours) 0.233
RNAGenScape (ours) 0.235
In addition to being successful in optimizing properties, RNAGenScape also achieves the best manifold fidelity, as shown in Table 2. See Appendix B for more details on the evaluations.
4.4 Efficiency and scalability
In addition to its strong property control, RNAGenScape is also highly efficient at inference time. As reported in Table 1, it achieves an inference speed of 0.57 ms/sample, nearly matching the fastest method (gradient ascent at 0.50 ms/sample) and substantially faster than many other property optimization methods (such as hill climbing at 81.53 ms/sample, DiffAb at 41.04 ms/sample, and IgLM at 157.57 ms/sample). This efficiency makes RNAGenScape well suited for large-scale or iterative design workflows where fast feedback is essential.
4.5 Interpolating between arbitrary sequences
RNAGenScape enables interpolation between arbitrary sequences using the directional drift term (equation 4). Guided by a directional force toward a specified target, RNAGenScape generates smooth and coherent trajectories on the learned manifold while preserving biological plausibility and continuity (Figure 3). These trajectories connect arbitrary input-target sequence pairs in a structured manner, reflecting semantically meaningful transitions.
The distances from each intermediate point to the source and target quantitatively demonstrate the monotonicity and smoothness of the interpolation (Figure 4).
Figure 3: Latent space interpolation trajectories from 5 sources to 4 targets. Each trajectory is shown as a line fading from bright to dark in a consistent color. RNAGenScape produces smooth and coherent paths on the manifold between arbitrary input-target mRNA pairs.
Figure 4: Latent space $\ell_{2}$ distances during interpolation show smooth and monotonic transition from the source to the target. Results are averaged over all data samples.
4.6 Ablation studies
Table 3: Manifold projector $\Psi$ is critical.
@@ 4.6 Ablation studies
latent space
distance $\downarrow$
$\Psi$ $\Delta$ $\uparrow$ % $\uparrow$ $\Delta$ $\downarrow$ % $\uparrow$
-0.17 45.1 -0.13 47.0 0.355
-0.46 72.4 -1.66 87.3 0.120
Manifold projector
Our first ablation shows that the manifold projector $\Psi$, a core contribution of RNAGenScape, is essential for its performance (Table 3). Without $\Psi$, the method completely fails at property optimization. This outcome is expected: following the property gradient without projecting back to the manifold causes trajectories to drift away, as reflected in the tripled latent space distance in Table 3.
Optimization steps
Next, we analyze how sensitive RNAGenScape is to the number of optimization steps. The results in Figure 5 show that our method is able to converge quickly and remains stable over a range of Langevin dynamics steps.
SUGAR
@@
5 Conclusion
We introduced RNAGenScape, a property-guided manifold Langevin dynamics framework dedicated for mRNA design and optimization. By combining an organized autoencoder that aligns the latent space with target properties and a denoising-based manifold projector, RNAGenScape steers existing sequences along smooth, manifold-aligned trajectories that both improve target properties and preserve biological plausibility. Empirically, RNAGenScape outperforms various generative and optimization methods in property control and manifold fidelity, while matching or exceeding their inference efficiency. It also shows promises in faithfully interpolating between real biological sequences. With this work, we also hope to shift the paradigm of biological sequence design from unconstrained generation to guided optimization, and shed more light on mRNA sequence design as a critical yet understudied frontier in computational biology.
6 Limitations and Future Work
One limitation of our approach is its dependence on the fidelity of the organized latent space: if the organized autoencoder fails to capture critical sequence constraints, manifold projections may permit small but functionally invalid drifts. Additionally, our current formulation optimizes a single scalar property; extending RNAGenScape to multi-objective settings would broaden its applicability. Finally, while we have demonstrated compelling in silico gains, integrating real-world experimental feedback remains an important avenue to validate and refine the learned manifold.
In future work, We will study the possibility to perform sequence-structure joint modeling and optimization. Beyond mRNA, we plan to extend RNAGenScape to other modalities such as protein sequences and regulatory elements, and integrate active learning frameworks that guide wet lab experimentation. By grounding sequence optimization in the manifold of real data, we aim to provide a versatile platform for interpretable and high-throughput design in synthetic biology.

### v2

@@ 4 Empirical Results
We evaluate RNAGenScape on property-optimized mRNA sequence generation across multiple biologically relevant settings. The goal is to improve target properties such as translation efficiency and stability while preserving the biological viability of the generated sequences. To this end, we conducted comprehensive experiments to analyze the property gains, biological viability, efficiency, and optimization trajectories.
Table 1: Our proposed RNAGenScape generates biologically viable mRNA sequences that achieve superior property optimization. In property optimization results, $\Delta$ denotes the median change in property and $\%$ denotes success rate (percentage of mRNAs improved). In the biological viability checks, uORF OOF AUG denote out-of-frame start codon in the upstream open reading frame, and Kozak similarity is the similarity to the consensus Kozak sequence. The best values per metric are highlighted in bold.
Method Property Optimization Biological Viability Checks
++Property -Property Initiation Interference Initiation Context Structural Stability
[…]
Data in the test set $1.65\pm 0.23$ $37.90\pm 30.38$ $30.13\pm 17.76$
DiffAb (Luo et al., 2022) +0.06+0.06 $55.0$ +0.11+0.11 $44.0$ $1.75\pm 2.29$ $27.79\pm 30.03$ $21.10\pm 12.46$
[…]
NOS-D (Gruver et al., 2023) +0.18+0.18 $58.3$ 0.29-0.29 $65.4$ $2.27\pm 2.52$ $32.40\pm 30.06$ $18.81\pm\phantom{0}5.82$
RNAGenScape (ours) $\textbf{0.54}$ 77.5 $\textbf{2.81}$ 97.9 $\textbf{1.63}\pm 0.59$ $\textbf{34.65}\pm 23.26$ $\textbf{25.78}\pm\phantom{0}3.17$
Gain over baselines $125\% +$$22.8\% +$$24.9\% +$$8.1\%$
[…]
Data in the test set $2.51\pm 2.65$ $33.92\pm 29.42$ $29.40\pm 7.20$
DiffAb (Luo et al., 2022) 0.21-0.21 $36.3$ 0.21-0.21 $60.8$ $3.75\pm 2.65$ $23.84\pm 29.41$ $27.31\pm 7.19$
[…]
RNAGenScape (ours) $\textbf{0.77}$ 75.0 $\textbf{1.29}$ 85.1 $\textbf{2.41}\pm 1.98$ $\textbf{26.17}\pm 17.43$ $\textbf{29.24}\pm 6.17$
Gain over baselines $148\% +$$30.4\% +$$20.6\% +$$5.3\%$
[…]
Data in the test set $3.32\pm 4.25$ $20.60\pm 29.54$ $8.65\pm 3.66$
DiffAb (Luo et al., 2022) 0.05-0.05 $45.0$ 0.04-0.04 $54.2$ $3.16\pm 4.19$ $14.68\pm 28.77$ $7.25\pm 3.57$
IgLM (Shuai et al., 2023) $\textbf{0.63}$ $80.5$ 0.51-0.51 $65.5$ $6.14\pm 4.85$ $\phantom{0}7.73\pm 20.51$ $7.27\pm 3.38$
NOS-C (Gruver et al., 2023) +0.08+0.08 $53.6$ +0.10+0.10 $46.0$ $4.53\pm 3.89$ $12.96\pm 22.88$ $5.79\pm 2.98$
NOS-D (Gruver et al., 2023) 0.09-0.09 $46.5$ 0.10-0.10 $53.6$ $5.64\pm 5.06$ $\phantom{0}8.05\pm 16.27$ $0.77\pm 1.35$
RNAGenScape (ours) $\textbf{0.63}$ 81.4 $\textbf{0.58}$ 67.8 $\textbf{4.51}\pm 3.68$ $\textbf{16.41}\pm 28.42$ $\textbf{8.23}\pm 3.18$
Gain over baselines $0.0\% +$$1.1\% +$$13.7\% +$$3.5\%$
@@ 4.1 Experimental settings
Datasets and the properties to optimize
We evaluated RNAGenScape on three mRNA datasets that capture diverse contexts and experimental designs. The optimization objectives are underlined.
@@
    OpenVaccine contains 2,400 mRNA sequences devised for COVID-19 mRNA vaccines, each with 107 nucleotides (Das et al., 2020). The target property is mRNA stability, which is directly relevant to vaccine performance.
@@
    Zebrafish consists of five subsets of zebrafish 5’ UTR mRNA sequences each with 124 nucleotides, comprising approximately 55,000 sequences in total. The target property is translation efficiency, which measures how effectively an mRNA sequence is translated into protein.
@@
    Ribosome is a large-scale dataset of approximately 260,000 5’ UTR mRNA sequences, each with 50 nucleotides (Sample et al., 2019). The target property is the mean ribosome load, defined as the average number of ribosomes bound to the mRNA during translation.
@@
All experiments were repeated using 5 random seeds and the average results are reported. Details are summarized in Supplementary Section A.
4.2 RNAGenScape generates biologically viable, property-optimized mRNA sequences
Across all three datasets, we evaluate RNAGenScape along two complementary axes: improvement of the target property and preservation of biological viability. Both aspects are jointly reported in Table 1.
Property optimization is assessed using UTR-LM (Chu et al., 2024), a widely adopted mRNA language model, as the external oracle. UTR-LM is first pre-trained on over 480,000 5’ UTR sequences from diverse sources and subsequently fine-tuned for property prediction on each dataset. After fine-tuning, the model is frozen and never accessed during optimization, ensuring an unbiased evaluation of the optimized sequences.
To assess biological viability, we apply a set of complementary checks grounded in established principles of mRNA translation. Specifically, we consider three classes of viability metrics. First, upstream out-of-frame start codons (uORF OOF AUG) are known to suppress downstream protein expression and are therefore undesirable (Dever et al., 2023). Second, Kozak similarity (Gleason et al., 2022) measures how closely the initiation context matches the consensus Kozak sequence, with higher values indicating more favorable translation initiation. Third, minimum free energy (Lorenz et al., 2011) serves as a proxy for mRNA structural stability.
Across all three datasets and both optimization directions, RNAGenScape achieves the largest median property change and the highest success rate (Table 1). Compared to state-of-the-art biological sequence optimization methods, including DiffAb (Luo et al., 2022), IgLM (Shuai et al., 2023), NOS-C (Gruver et al., 2023), and NOS-D (Gruver et al., 2023), RNAGenScape attains up to a 148% increase in median property improvement and up to a 30% increase in success rate. In addition to better optimization performance, RNAGenScape consistently produces sequences with fewer aberrant upstream start codons, higher Kozak similarity, and more favorable minimum free energy profiles. Moreover, the viability metrics of RNAGenScape-generated sequences more closely match those of real sequences in the unseen test set, indicating that substantial property gains can be achieved without compromising biological viability.
4.3 Held-out validation on unseen near-optimal sequences
Table 2: Edit distance between each generated sequence and its nearest held-out sequence with near-optimal property. Edit distance is normalized to $[0,1]$, with lower values being better.
DiffAb IgLM NOS-C NOS-D RNAGenScape
(Luo et al., 2022) (Shuai et al., 2023) (Gruver et al., 2023) (Gruver et al., 2023) (ours)
@@ 4.3 Held-out validation on unseen near-optimal sequences
To further assess whether RNAGenScape produces realistic property-optimized solutions, we conducted a held-out validation designed to mimic a practical design scenario. Specifically, for each dataset, we withheld a subset of sequences that already exhibit near-optimal target properties and excluded them from optimization. The property threshold is set at one standard deviation better than the mean. We then evaluate how close the generated sequences are to these held-out high-quality references.
Table 2 reports the normalized edit distance between each generated sequence and its nearest held-out sequence with a near-optimal property. Across all datasets, RNAGenScape consistently achieves the lowest edit distance compared to alternative biological sequence generation methods, indicating that it converges to realistic, high-performing sequences. These results suggest that RNAGenScape not only improves target properties, but does so in a manner that recovers solutions closely aligned with experimentally observed sequences.
4.4 Efficiency and scalability
Figure 2: Inference throughput of all optimization methods.
In addition to its other merits, RNAGenScape is also highly efficient at inference time. As reported in Figure 2, it achieves a throughput of 1.754 sample/ms, 68% faster than the most efficient biological sequence optimization method being compared. This efficiency makes RNAGenScape well suited for large-scale or iterative design workflows where high throughput is essential.
4.5 Ablation studies
Table 3: Reconstruction and prediction objectives in organized autoencoder are constructive and complementary.
$\mathcal{L}_{\text{Recon}}$ $\mathcal{L}_{\text{Pred}}$ Reconstruction Property Prediction
Nucleotide Acc. $\uparrow$ Pearson Corr. $\uparrow$ Spearman Corr. $\uparrow$
@@ 4.5 Ablation studies
0.17  0.75  0.75
0.67  0.70  0.71
@@
0.17  0.54  0.60
0.78  0.63  0.67
[…]
0.11  0.87  0.85
0.99  0.87  0.85
Table 4: The manifold projector $\Psi$ substantially improves property optimization by enforcing optimization along the data manifold.
$\Psi$ ++property -property
$\Delta$ $\uparrow$ % $\uparrow$ $\Delta$ $\downarrow$ % $\uparrow$
[…]
   0.54  77.5  $\textbf{2.81}$  97.9
[…]
   0.77  75.0  $\textbf{1.29}$  85.1
[…]
   0.63  81.4 $\textbf{0.58}$  67.8
Organized autoencoder
Since the optimization process in RNAGenScape is performed in the learned latent manifold, we verify that the OAE provides both accurate reconstruction and property prediction (Table 3). Additionally, we demonstrate that the reconstruction and prediction objectives ($\mathcal{L}_{\text{Recon}}$ and $\mathcal{L}_{\text{Pred}}$) are constructive to each other, which justifies our OAE design.
Manifold projector
Next, we show that the manifold projector $\Psi$ is essential for the optimization performance of RNAGenScape (Table 4). Without $\Psi$, optimization becomes substantially less effective across datasets. This behavior is expected: updating latent representations solely by following the property gradient allows trajectories to drift away from the learned data manifold, entering regions that are weakly supported by real sequences. In these off-manifold regions, property gradients become less reliable, leading to unstable updates and diminished optimization gains.
Optimization steps
We further analyze how sensitive RNAGenScape is to the number of optimization steps. The results in Figure 3 show that our method is able to converge quickly and remains stable over a range of optimization steps during Langevin dynamics updates.
Table 5: RNAGenScape achieves better property optimization than de novo generative methods and classic optimization algorithms that operate on our learned data manifold.
Method Property Optimization
@@
$\Delta$ $\uparrow$ % $\uparrow$ $\Delta$ $\downarrow$ % $\uparrow$
@@
VAE (Kingma et al., 2013) 0.23-0.23 $42.1$ 0.23-0.23 $57.9$
DDPM (Ho et al., 2020) 0.33-0.33 $33.8$ 0.33-0.33 $66.2$
LDM (Rombach et al., 2022) 0.07-0.07 $47.5$ 0.07-0.07 $52.5$
FM (Lipman et al., 2022) 0.34-0.34 $32.5$ 0.34-0.34 $67.5$
OAE + gradient ascent 0.01-0.01 $51.0$ 0.19-0.19 $61.3$
OAE + MCMC 0.11-0.11 $45.1$ 0.19-0.19 $57.7$
OAE + hill climbing +0.40+0.40 $69.2$ 0.29-0.29 $64.2$
OAE + stochastic hill climbing 0.12-0.12 $43.5$ 0.15-0.15 $60.0$
RNAGenScape (ours) $\textbf{0.54}$ 77.5 $\textbf{2.81}$ 97.9
[…]
VAE (Kingma et al., 2013) 0.01-0.01 $49.4$ 0.01-0.01 $50.6$
DDPM (Ho et al., 2020) +0.29+0.29 $58.2$ +0.29+0.29 $41.8$
LDM (Rombach et al., 2022) 0.95-0.95 $22.5$ 0.95-0.95 $77.5$
FM (Lipman et al., 2022) +0.32+0.32 $59.8$ +0.32+0.32 $40.2$
OAE + gradient ascent 0.21-0.21 $18.0$ 0.33-0.33 $66.5$
OAE + MCMC 0.20-0.20 $44.3$ 0.37-0.37 $60.1$
OAE + hill climbing +0.76+0.76 $74.4$ 0.87-0.87 $75.2$
OAE + stochastic hill climbing +0.32+0.32 $60.3$ 0.80-0.80 $73.7$
RNAGenScape (ours) $\textbf{0.77}$ 75.0 $\textbf{1.29}$ 85.1
[…]
VAE (Kingma et al., 2013) 0.24-0.24 $41.7$ 0.24-0.24 $58.3$
DDPM (Ho et al., 2020) 0.11-0.11 $45.9$ 0.11-0.11 $54.1$
LDM (Rombach et al., 2022) 0.24-0.24 $41.8$ 0.24-0.24 $58.2$
FM (Lipman et al., 2022) 0.10-0.10 $46.4$ 0.10-0.10 $53.6$
OAE + gradient ascent +0.43+0.43 $73.6$ 0.05-0.05 $55.7$
OAE + MCMC 0.19-0.19 $43.0$ 0.10-0.10 $54.2$
OAE + hill climbing +0.53+0.53 83.0 0.06-0.06 $56.1$
OAE + stochastic hill climbing +0.19+0.19 $60.4$ 0.05-0.05 $55.2$
RNAGenScape (ours) $\textbf{0.63}$ $81.4$ $\textbf{0.58}$ 67.8
De novo generative methods and classic optimization algorithms
In addition to state-of-the-art biological sequence optimization methods, we compare RNAGenScape with two different types of methods. The first type is de novo generative modeling approaches, namely variational autoencoder (VAE) (Kingma et al., 2013), denoising diffusion probabilistic model (DDPM) (Ho et al., 2020), latent diffusion model (LDM) (Rombach et al., 2022), and flow matching (FM) (Lipman et al., 2022). The second type are classic optimization algorithms that operate in our OAE latent manifold, including gradient ascent (Williams, 1992; Zinkevich, 2003), Markov chain Monte Carlo (MCMC) (Brooks, 1998; Andrieu et al., 2003), and hill climbing (Selman and Gomes, 2006). All classic optimization baselines are GPU-compatible re-implementations of Castro et al. (Castro et al., 2022).
The results in Table 5 show that 1) de novo generative models exhibit limited property control and 2) classic optimization strategies in the OAE latent manifold also underperform RNAGenScape, indicating that our manifold Langevin dynamics contributed substantially to the favorable performance. See Figure S1 for the optimization performance of all methods being compared.
4.6 Case study: RNAGenScape produces structured, data-aligned optimization trajectories
Figure 4: A sample latent space trajectory of RNAGenScape over 10 optimization steps. The trajectory follows a smooth path with steady improvement in the target property. a, The trajectory in the PHATE space and the corresponding 3D structures. b, Intermediate results of RNAGenScape and competing methods midway through property optimization (at the 5th step out of 10 steps).
To illustrate how RNAGenScape iteratively navigates along a learned latent manifold, we visualize a sample optimization run in Figure 4a. The trajectory exhibits monotonic increases in the target property, while remaining near regions populated by real sequences. See Supplementary Section E for more examples. These trajectories are direct consequences of the manifold-constrained dynamics, which guided each step toward high-property regions while staying on the manifold.
As a qualitative illustration of biological plausibility, we show that the intermediate results of RNAGenScape optimization can be properly folded by RhoFold (Shen et al., 2024) into 3D structures (Figure 4b). In contrast, folding intermediate products of several other methods causes failure, as indicated by fragmented structures (Figure 4b).